OpenStudy (anonymous):

The population (in millions) of a certain island can be approximated by the function: P(x)=50 (1.05)^x Where x is the number of years since 2000. In which year will the population reach 200 million? Hint: An answer such as 2002.4 would represent the year 2002. A) 2028 B)2066 C) 2002 D) 2015

OpenStudy (anonymous):

@freckles

OpenStudy (danjs):

solve 200 = 50*1.05^x x is in millions of years

OpenStudy (danjs):

do you know how to begin?

OpenStudy (anonymous):

no...

OpenStudy (danjs):

\[200 = 50*1.05^x\]

OpenStudy (danjs):

divide both sides by 50 first

OpenStudy (danjs):

\[4 = 1.05^x\]

OpenStudy (danjs):

recall: \[\ln(b^x) = x*\ln(b)\]

OpenStudy (danjs):

take ln of both sides

OpenStudy (anonymous):

the whole ln thing confuses me. do i put 1.05 and ln into the calculator?

OpenStudy (danjs):

after dividing by 50 on both sides you are left with \[4 = 1.05^x\] You then take the natural log of both sides

OpenStudy (danjs):

\[\ln(4) = \ln(1.05^x)\] \[\ln(4) = x*\ln(1.05)\]

OpenStudy (danjs):

divide both sides now by ln(1.05) to isolate x by itself

OpenStudy (anonymous):

1.3862943611 = x(0.0487901642) Is this right?

OpenStudy (danjs):

\[x = \frac{ \ln(4) }{ \ln(1.05) }\]

OpenStudy (anonymous):

I mean this: 1.3862943611 / (0.0487901642) = x Then I just divide like normal right?

OpenStudy (danjs):

natural log of 4, divided by natural log of 1.05, idk i dont have a calculator on me

OpenStudy (danjs):

x should be about 28

OpenStudy (anonymous):

hold on...using my calculator

OpenStudy (danjs):

Where x is the number of years since 2000. In which year will the population reach 200 million? The final answer is: 2000 + (the x you calculated)

OpenStudy (anonymous):

Okay, I got roughly 2,002.8

OpenStudy (danjs):

i just got x = 28.413 about... so that rounds to 28

OpenStudy (danjs):

2000 + x = 2000 + 28 the year is 2028 when the population reaches 200 million

OpenStudy (danjs):

i hope you understood how logarithms were used to solve for x. Basically the property \[\ln(b^x) = x*\ln(b)\]

OpenStudy (anonymous):

Yeah, I got it. Thanks!

OpenStudy (danjs):

k, welcome